I would appreciate the kindness of your vauled help.
I have many to do!!!
Solve and show work!
(2y + 6) (y - 3) = 0

Answer by longjonsilver(2297)

if (2y + 6) (y - 3) = 0 then we can say that

either 2y+6=0 or y-3=0.

in which case, we have
2y=-6 or y=3
so y=-6/2 or y=3

hence y=-3 or y=3

jon.